OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2.

OpenStudy (anonymous):

I did it my way and it came out 10. Not sure if im doing it right tho

OpenStudy (ranga):

10 is correct for the third term. You still need to find the 4th and then the fifth term.

OpenStudy (anonymous):

Ok then im doing something wrong. can you explain to me how to do the 4th one correctly and then i can do the 5th one?

OpenStudy (ranga):

Sure. The first thing I would do is to simplify the original equation because you have two constants in the nth term. f(n) = f(1) + f(2) + f(n - 1) = 2 + 4 + f(n - 1) f(n) = 6 + f(n - 1) put n = 3 f(3) = 6 + f(2) = 6 + 4 = 10 put n = 4 f(4) = 6 + f(3) = 6 + 10 = 16

OpenStudy (anonymous):

Ok so f(5) will go like this? f(5) = 6 + f(4) and f(4) equaled 16 so f(5) = 6 + 16 = 22. f(5) = 22?

OpenStudy (ranga):

Yes!

OpenStudy (anonymous):

Just another one for practice. For f(6) f(6) = 6 + f(5) and f(5) was 22 so f(6) = 6 + 22 =28?

OpenStudy (ranga):

If you examine the f(n) formula you can quickly figure out how to easily get other terms. f(n) = 6 + f(n - 1) It says the nth term is 6 added to the previous term. So 6th term = 6 + 5th term = 6 + 22 = 28 7th term = 6 + 6th term = 34 Yes your sixth term is correct too. You got it!

OpenStudy (anonymous):

Thanks!

OpenStudy (ranga):

you are welcome.